OpenStudy (anonymous):

cos2a= cos^2 a- sin^2 a for all values of a. True or False

OpenStudy (anonymous):

That looks like an identity.

OpenStudy (ash2326):

\[\cos 2a=\cos^2 a-\sin ^2 a\] Yes, it's an identity. True

OpenStudy (hba):

Please prove it @ash2326

OpenStudy (ash2326):

We know that \[\cos (A+B)=\cos A \cos B-\sin A \sin B\] here \[A=B=a\] so \[\cos (a+a)=\cos a \cos a -\sin a \sin a\] \[\cos 2a=\cos^2 a-\sin^2 a\]

OpenStudy (anonymous):

Further proof (using geometry) can be found here: http://www.themathpage.com/atrig/sum-proof.htm

OpenStudy (anonymous):

You can also prove it using complex numbers.